OpenStudy (anonymous):

Explain how -3/4 - 1 = -7/4

OpenStudy (anonymous):

-3/4-1=(-3-4)/4=-7/4

OpenStudy (anonymous):

Oh! to you turn 1 into 4/4 in order to have a common denominator 3and 4 7.. thanks

hero (hero):

-3/4 - 1 = -7/4 7/4 - 3/4 = 1 4/4 = 1 1 = 1 True

OpenStudy (anonymous):

yup..